# Anomalous dynamical scaling in interacting anyonic chains

### 16 Mar 2026 v1

@@ Abstract
In the main text, we presented a representative example of the propagation of correlations from $d=1$ to $d=3$, corresponding to the leftward propagation of the holon. Here we provide the complementary process associated with the rightward propagation of the doublon, in Fig. S6(a). This contribution also involves two interfering paths: (i) two adjacent particles tunnel sequentially to the right (orange arrows), exchanging once and yielding an amplitude $2J^{2}e^{-i\theta}$; (ii) the left particle tunnels twice (green arrows), exchanging twice with its neighbor and acquiring an anyonic phase $4J^{2}e^{-i2\theta}$. The total amplitude arises from the coherent sum of these processes. In the bosonic limit ($\theta=0$), the two paths interfere constructively, enhancing the propagation of the holon-doublon pair, whereas in the fermionic limit ($\theta=\pi$), the interference becomes destructive, strongly suppressing the propagation.
@@
This Floquet approach provides an effective route to realize density-dependent hopping phases, as illustrated in Fig. S7(b). Defining the effective on-site interaction as $U$, the primary processes are: (i) $|1,0\rangle\to|0,1\rangle$, driven at frequency $\omega_{1}=E$; (ii) $|1,1\rangle\to|0,2\rangle$, driven at $\omega_{2}=E-(U_{0}-U)$; (iii) $|2,0\rangle\to|1,1\rangle$, driven at $\omega_{3}=E+(U_{0}+U)$. To map these processes onto the anyon Hubbard model, an additional phase offset $\theta$ is imprinted on the drive for process (iii), while processes (i) and (ii) remain reference transitions [gray block in Fig. S7(b)]. This configuration is sufficient for low fillings; however, at unit filling, an additional process (iv) $|2,1\rangle\to|1,2\rangle$ emerges as an important additional contribution. Due to the energy offset $E$, this transition is unavoidably induced by the same Floquet channel as process (i) and lacks the required phase (orange block).

### 22 Mar 2026 v2

@@ Abstract
In the main text, we presented a representative example of the propagation of correlations from $d=1$ to $d=3$, corresponding to the leftward propagation of the holon. Here we provide the complementary process associated with the rightward propagation of the doublon in Fig. S6(a). This contribution also involves two interfering paths: (i) two adjacent particles tunnel sequentially to the right (orange arrows), exchanging once and yielding an amplitude $2J^{2}e^{-i\theta}$; (ii) the left particle tunnels twice (green arrows), exchanging twice with its neighbor and acquiring an anyonic phase $4J^{2}e^{-i2\theta}$. The total amplitude arises from the coherent sum of these processes. In the bosonic limit ($\theta=0$), the two paths interfere constructively, enhancing the propagation of the holon-doublon pair, whereas in the fermionic limit ($\theta=\pi$), the interference becomes destructive, strongly suppressing the propagation.
@@
This Floquet approach provides an effective route to realize density-dependent hopping phases, as illustrated in Fig. S7(b). The Hamiltonian is
$\hat{H}^{\prime}(t)=$ $-J\sum_{j=1}^{L-1}\left(\hat{b}_{j}^{\dagger}\hat{b}_{j+1}+\text{h.c.}\right)+\frac{U_{0}}{2}\sum_{j=1}^{L}\hat{n}_{j}(\hat{n}_{j}-1)+E\sum_{j=1}^{L}j\hat{n}_{j}+\sum_{i}V_{i}(t).$ (S22)
Defining the effective onsite interaction as $U$, the primary processes are: (i) $|1,0\rangle\to|0,1\rangle$, driven at frequency $\omega_{1}=E$; (ii) $|1,1\rangle\to|0,2\rangle$, driven at $\omega_{2}=E-(U_{0}-U)$; (iii) $|2,0\rangle\to|1,1\rangle$, driven at $\omega_{3}=E+(U_{0}+U)$. To map these processes onto the anyon Hubbard model, an additional phase offset $\theta$ is imprinted on the drive for process (iii), while processes (i) and (ii) remain reference transitions [gray block in Fig. S7(b)]. This configuration is sufficient for low fillings; however, at unit filling, an additional process (iv) $|2,1\rangle\to|1,2\rangle$ emerges as an important additional contribution. Due to the energy offset $E$, this transition is unavoidably induced by the same Floquet channel as process (i) and lacks the required phase (orange block).